# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 5.3 Error estimates
For the estimate (5.37a), let us make a case distinction for the terms in this sum, depending on $\ell_{1}$. For $20<\ell_{1}\leq N-$ 2 we have
$e^{\ell Hs}\|\nabla^{\ell_{1}}(\Phi k)\star\nabla^{\ell_{2}}\mathcal{T}\|_{L^{2}}\leq e^{\ell_{1}Hs}\|\nabla^{(\ell_{1})}(\Phi k)\|_{L^{\infty}}\|\mathcal{T}\|_{H^{\ell_{2}}}\leq Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell-1}}$ (5.38)
@@
$\nabla^{(\ell_{1})}\bigl{(}\Phi k\bigr{)}=\nabla^{(\ell_{1})}\bigl{(}\widehat{\Phi}\widehat{k}+\widetilde{\Phi}\widehat{k}+\widehat{\Phi}\widetilde{k}+\widetilde{\Phi}\widetilde{k}\bigr{)}$ (5.39)
involving at most $N-2$ derivatives of $\widehat{\Phi}$, and $\widehat{k}$. For the remaining terms, with more that $N-2$ derivatives of $\widehat{\Phi}$ or $\widehat{k}$, we use the bootstrap assumption (4.12) on the energy, and Lemma 4.14 for $\mathcal{T}$. For example:
@@
$\begin{split}\Phi\widehat{\nabla}_{i}\widetilde{k}_{c}{}^{a}&+\Phi\widehat{\nabla}_{c}\widetilde{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\widehat{\nabla}_{b}\widetilde{k}_{i}{}^{j}=\\ =&\,\Phi\widehat{\Gamma}_{ib}^{a}\widetilde{k}_{c}{}^{b}-\Phi\widehat{\Gamma}_{ic}^{b}\widetilde{k}_{b}{}^{a}+\Phi\widehat{\Gamma}_{cb}^{a}\widetilde{k}_{i}{}^{b}-\Phi\widehat{\Gamma}_{ic}^{b}\widetilde{k}_{b}{}^{a}-g^{ab}g_{cj}\Phi(\widehat{\Gamma}_{bc}^{j}\widetilde{k}_{i}{}^{c}-\widehat{\Gamma}_{bi}^{c}\widetilde{k}_{c}{}^{j})\\ =&\,\Phi\widehat{\Gamma}_{ib}^{a}(\widetilde{k}_{c}{}^{b}-H\delta_{c}{}^{b})-\Phi\widehat{\Gamma}_{ic}^{b}(\widetilde{k}_{b}{}^{a}-H\delta_{b}{}^{a})+\Phi\widehat{\Gamma}_{cb}^{a}(\widetilde{k}_{i}{}^{b}-H\delta_{i}{}^{b})-\Phi\widehat{\Gamma}_{ic}^{b}(\widetilde{k}_{b}{}^{a}-H\delta_{b}{}^{a})\\ &-g^{ab}g_{cj}\Phi\big{[}\widehat{\Gamma}_{bc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})-\widehat{\Gamma}_{bi}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})\big{]}\end{split}$ (5.44)
@@
$\begin{split}k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi}=&\widehat{k}_{c}{}^{a}\nabla_{i}\widehat{\Phi}+\widehat{k}_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}\widehat{k}_{i}{}^{j}\nabla_{b}\widehat{\Phi}\\ +(\widetilde{k}_{c}{}^{a}-H\delta_{c}{}^{a})\nabla_{i}\widehat{\Phi}&+(\widetilde{k}_{i}{}^{a}-H\delta_{i}{}^{a})\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})\nabla_{b}\widehat{\Phi}\\ &+\delta_{c}{}^{a}\nabla_{i}\widehat{\Phi}+\delta_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{ci}\nabla_{b}\widehat{\Phi}\end{split}$ (5.45)
@@
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}(k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi})\|_{L^{2}(\Sigma_{s},g)}\\ \leq\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{-\frac{1}{2}Hs}e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}$ (5.46)
@@
$\mathrm{Error}_{\widehat{g},\ell}=$ $\nabla^{(\ell)}\mathrm{Error}_{\widehat{g}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}$ (5.52)
$(\mathrm{Error}_{\widehat{g}})_{ij}=$ $2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}$ (5.53)
@@
For the terms involving derivatives of the reference metric, namely $\nabla^{(\ell)}\widetilde{\Phi}$ or $\nabla^{(\ell)}\widetilde{g}$, we also separate differences using $\nabla=\widehat{\nabla}+\widetilde{\nabla}$ thus introducing $\widehat{\Gamma}$ as a quantity that can be treated alongside $\widehat{k}$ as above. The remaining terms ~$\widetilde{\nabla}^{(\ell)}\widetilde{\Phi}$ and ~$\widetilde{\nabla}^{(\ell)}\widetilde{g}$ can always be estimated in $L^{\infty}$. For example,
@@ 5.4 Main energy estimates
$\Phi k-H\delta=\widehat{\Phi}\widehat{k}+\widetilde{\Phi}\widehat{k}+\widehat{\Phi}\widetilde{k}+\widetilde{\Phi}\widetilde{k}-H\delta$ (5.61)
@@
$\partial_{s}(e^{-3Hs}\mathrm{vol}_{g})=(\Phi\,\text{tr}_{g}k-3H)e^{-3Hs}\mathrm{vol}_{g},\qquad|\Phi\,\text{tr}_{g}k-3H|\leq Ce^{-Hs},$ (5.63)
@@
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}\nabla^{i}\big{[}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big{]}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\\ =-\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}g^{ii^{\prime}}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big{[}\nabla_{i^{\prime}}f^{2}(t)\big{]}e^{-3Hs}\mathrm{vol}_{g}\\ \leq\,Ce^{-Hs}\mathcal{E}_{N}(s)+\frac{1}{2}e^{-Hs}e^{3Hs}e^{2NHs}\|\nabla^{(N+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}^{2}$ (5.66)
@@ 6 Precise asymptotics of the perturbed solution
$\begin{split}\widehat{g}_{ij}(s,x)=\widehat{g}_{ij}^{\infty}(x)e^{2Hs}+\widehat{h}_{ij}(s,x),\\ \widehat{\Phi}(s,x)=\widehat{\Phi}^{\infty}(x)e^{-2Hs}+\widehat{\Psi}(s,x)\,,\end{split}$ (6.2)
@@
$\begin{split}\|f(t)\widehat{g}^{\infty}(x)\|_{W^{N-4}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\qquad\|f(t)\widehat{h}(s,x)\|_{W^{N-4}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\\ \|f(t)\widehat{\Phi}^{\infty}(x)\|_{W^{N-6}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\qquad\|f(t)\widehat{\Psi}(s,x)\|_{W^{N-6}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-4Hs}\,,\end{split}$ (6.3)
@@
$\Rightarrow\quad\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq$ $\,e^{-3Hs}e^{3Hs_{0}}\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s_{0}},g)}+C\mathring{\varepsilon}e^{-\frac{3}{2}Hs}\leq C\mathring{\varepsilon}e^{-\frac{3}{2}Hs},$
@@
Repeating the above argument, integrating in $[s_{0},s]$, gives
@@
Integrating in $[s_{0},s]$ and repeating the above argument gives
@@
Next, we employ the already derived (6.1), together with (4.19), to estimate the RHS of (2.37):
@@
$\bigg{\|}\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{-2Hs}\widehat{g})\mathrm{d}s\bigg{\|}_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\qquad s_{1}<s_{2}.$
@@
On the other hand, integrating in $[s,+\infty)$ gives:
@@
[…]
@@
$\bigg{\|}\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{2Hs}\widehat{\Phi})\mathrm{d}s\bigg{\|}_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\qquad s_{1}<s_{2}.$
Hence, $e^{2Hs}\widehat{\Phi}$ has a $W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ limit, as $s\rightarrow+\infty$, denoted by $\widehat{\Phi}^{\infty}(x)$. Moreover, with $\widehat{\Psi}=\widehat{\Phi}-\widehat{\Phi}^{\infty}e^{-2Hs}$, integrating in $[s,+\infty)$ gives:

### v2

@@ 5.3 Error estimates
For the estimate (5.37a), let us make a case distinction for the terms in this sum, depending on $\ell_{1}$. For $0<\ell_{1}\leq N-2$ we have
$e^{\ell Hs}\|\nabla^{\ell_{1}}(\Phi k)\star\nabla^{\ell_{2}}\mathcal{T}\|_{L^{2}}\leq e^{\ell_{1}Hs}\|\nabla^{(\ell_{1})}(\Phi k)\|_{L^{\infty}}\|\mathcal{T}\|_{H^{\ell_{2}}}\leq Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell-1}}\,$, (5.38)
@@
[…]
involving at most N2N-2 derivatives of $\widehat{\Phi}$, and $\widehat{k}$. For the remaining terms, with more that N2N-2 derivatives of $\widehat{\Phi}$ or $\widehat{k}$, we use the bootstrap assumption (4.12) on the energy, and Lemma 4.14 for $\mathcal{T}$. For example:
@@
[…]
@@
[…]
@@
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}(k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi})\|_{L^{2}(\Sigma_{s},g)}\\ \leq\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{-\frac{1}{2}Hs}e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}\,.$ (5.46)
@@
$\mathrm{Error}_{\widehat{g},\ell}$= $\nabla^{(\ell)}\mathrm{Error}_{\widehat{g}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}\,$, (5.52)
$(\mathrm{Error}_{\widehat{g}})_{ij}=$ $2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}\,.$ (5.53)
@@
For the terms involving derivatives of the reference metric, namely $\nabla^{(\ell)}\widetilde{\Phi}$ or $\nabla^{(\ell)}\widetilde{g}$, we also separate differences using $\nabla=\widehat{\nabla}$$\widetilde{\nabla}$ thus introducing $\widehat{\Gamma}$ as a quantity that can be treated alongside $\widehat{k}$ as above. The remaining terms $\widetilde{\nabla}^{(\ell)}\widetilde{\Phi}$ and $\widetilde{\nabla}^{(\ell)}\widetilde{g}$ can always be estimated in $L^{\infty}$. For example,
@@ 5.4 Main energy estimates
$\Phi k-H\delta=\widehat{\Phi}\widehat{k}+\widetilde{\Phi}\widehat{k}+\widehat{\Phi}\widetilde{k}+\widetilde{\Phi}\widetilde{k}-H\delta\,$, (5.61)
@@
$\partial_{s}(e^{-3Hs}\mathrm{vol}_{g})=(\Phi\,\text{tr}_{g}k-3H)e^{-3Hs}\mathrm{vol}_{g},\hskip 18.49988pt|\Phi\,\text{tr}_{g}k-3H|\leq Ce^{-Hs},$ (5.63)
@@
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}\nabla^{i}\big[(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big]f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\\ =-\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}g^{ii^{\prime}}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big[\nabla_{i^{\prime}}f^{2}(t)\big]e^{-3Hs}\mathrm{vol}_{g}\\ \leq\,Ce^{-Hs}\mathcal{E}_{N}(s)+\frac{1}{2}e^{-Hs}e^{3Hs}e^{2NHs}\|\nabla^{(N+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}^{2}\,$, (5.66)
@@ 6 Precise asymptotics of the perturbed solution
[…]
@@
[…]
@@
$\Rightarrow\qquad\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq$ $\,e^{-3Hs}e^{3Hs_{0}}\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s_{0}},g)}+C\mathring{\varepsilon}e^{-\frac{3}{2}Hs}\leq C\mathring{\varepsilon}e^{-\frac{3}{2}Hs},$
@@
Repeating the above argument, integrating in $s_{0},s]$, gives
@@
Integrating in $s_{0},s]$ and repeating the above argument gives
@@
Next, we employ the already derived (6.1), together with (4.19), to estimate the RHS of (2.35a):
@@
$\bigg\|\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{-2Hs}\widehat{g})\mathrm{d}s\bigg\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\hskip 18.49988pts_{1}<s_{2}.$
@@
On the other hand, integrating in $s,+\infty)$ gives:
@@
$\Rightarrow\qquad$ $\left\{\begin{array}[]{ll}\|f(t)e^{-2Hs}\widehat{h}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs}\\ \|f(t)\widehat{g}^{\infty}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq e^{-2Hs}\|f(t)\widehat{g}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}+C\mathring{\varepsilon}e^{-2Hs}\end{array}\right$.
$\Rightarrow\qquad$ $\left\{\begin{array}[]{ll}\|f(t)\widehat{h}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}\\ \|f(t)\widehat{g}^{\infty}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}\,.\end{array}\right$.
@@
$\bigg\|\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{2Hs}\widehat{\Phi})\mathrm{d}s\bigg\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\hskip 18.49988pts_{1}<s_{2}.$
Hence, $e^{2Hs}\widehat{\Phi}$ has a $W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ limit, as $s\rightarrow+\infty$, denoted by $\widehat{\Phi}^{\infty}(x)$. Moreover, with $\widehat{\Psi}=\widehat{\Phi}-\widehat{\Phi}^{\infty}e^{-2Hs}$, integrating in $s,+\infty)$ gives:
@@
$\Rightarrow\qquad$ $\left\{\begin{array}[]{ll}\|f(t)e^{2Hs}\widehat{\Psi}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs}\\ \|f(t)\widehat{\Phi}^{\infty}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq e^{2Hs}\|f(t)\widehat{\Phi}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}+C\mathring{\varepsilon}e^{-2Hs}\end{array}\right$.
$\Rightarrow\qquad$ $\left\{\begin{array}[]{ll}\|f(t)\widehat{\Psi}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-4Hs}\\ \|f(t)\widehat{\Phi}^{\infty}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}\,.\end{array}\right$.
This completes the proof of the proposition. ∎
Remark 6.2 (Asymptotic expansion).
Higher order terms in the asymptotic expansion can be obtained — assuming successively more derivatives for the solution — by plugging (6.2) into the equations (2.35a), (2.42) and repeating the arguments in the proof of the previous proposition. Similarly, we can derive asymptotic expansions for $\widehat{\Gamma}_{ic}^{a},\widehat{k}_{i}{}^{j}$; see for example [Fou22, Theorem 1.3]. A complete asymptotic expansion, and smoothness of the asymptotically rescaled metric across the conformal boundary at infinity $\mathcal{I}^{+}=\{s=\infty\}$ has recently been obtained in [HV24].
Appendix A Local well-posedness
For completeness, we outline a proof of the local well-posedness of the Einstein vaccum equations with cosmological constant in parabolic gauge.
The argument for local existence is qualitatively different from the proof of global existence for the following reason: In the derivation of the energy identity (5.17) for the variables $\widehat{k},\widehat{\Gamma}$ (which lies at the heart of the global existence argument), we use the momentum constraint (5.15), (5.16); see discussion in Section 5.1. Unfortunately, such an identity is not available in a contraction mapping argument (with the aim of proving local well-posedness), since the constraint equations are not valid at each step of an iteration scheme. A remedy for this problem is to consider a system of wave equations for $k_{ij}$, instead of the propagation equation (2.7) in the first place.
For the resulting reduced system of equations, standard results for quasilinear wave type systems can be used to establish local well-posedness. After showing that the solution produced is actually a solution to the Einstein vacuum equations (1.1), we infer as a corollary the local well-posedness of the original evolution equations in parabolic gauge (Section 2.2), i.e. existence, uniqueness and continuous dependence of solutions to initial data.
This type of argument was already employed in [CK93, Section 10.2.2] to prove the local well-posedness of the first and second variation equations (2.22a), (2.23), in the maximal gauge, where the corresponding equation for the lapse is elliptic:
$\operatorname{tr}k=0\qquad\triangle\Phi=|k|^{2}\Phi\,$,
see also [FS19] for the extension to an initial-boundary value problem. Similarly this argument was adapted in [RS18b, Section 14] for the CMC gauge $\operatorname{tr}k=f(t)$.
Now consider the equations (2.24c), (2.25) for $k_{i}{}^{j},\Gamma_{ic}^{a}$ which follow from (2.22a), (2.23). The parabolic gauge (2.31) can be written in the form $\operatorname{tr}k=\Phi+\text{a known source term}$. Since it makes no difference for local well-posedness, we drop the source term and consider for convenience the gauge condition:
$\Phi=\operatorname{tr}k.$ (A.1)
First we derive a second order (in time) equation satisfied by $k_{ij}$. Here the Einstein equations are not imposed. Instead we derive a geometric condition, adapted to (A.1) which is equivalent to a wave equation for $k_{ij}$.
Proposition A.1.
Let $s$ be a time function for the spacetime metric $\bm{g}$}, and $g,k$ the first and second fundamental forms as in (2.1) and (2.4). Then $k_{ij}$ satisfies the wave equation
$e_{0}^{2}k_{ij}-\Delta_{g}k_{ij}=I_{ij}+H_{ij}\,$, (A.2)
where
$H_{ij}=$ $-\Phi^{-3}\partial_{s}\Phi\nabla_{i}\nabla_{j}\Phi+\Phi^{-2}\nabla_{i}\nabla_{j}\partial_{s}\Phi-\nabla_{i}\nabla_{j}\Phi-\Phi^{-2}\partial_{s}\Gamma^{l}_{ij}\partial_{l}\Phi-e_{0}(k_{ij}\mathrm{tr}k-2{k_{i}}^{l}k_{jl})$ (A.3a)
$+\Phi^{-1}k_{ij}\Delta_{g}\Phi-\Phi^{-1}k_{i}{}^{a}\nabla_{a}\nabla_{j}\Phi-\Phi^{-1}k_{j}{}^{a}\nabla_{a}\nabla_{i}\Phi$
$-\Phi^{-1}\nabla^{a}\Phi(\nabla_{j}k_{ia}+\nabla_{i}k_{ja}-2\nabla_{a}k_{ij})+\Phi^{-1}\mathrm{tr}k\nabla_{j}\nabla_{i}\Phi$